OpenStudy (anonymous):

Which of the following is(are) equivalent to 3333 base seven + 3333 base seven? (i) 6666 base seven (ii) 1480 base twelve (iii) 10000 base seven A) (i) only B) (iii) only C) (i) and (ii) only D) (ii) and (iii) only

OpenStudy (anonymous):

these problems are all over the map what kind of class is this?

OpenStudy (anonymous):

Well we know (i) and (iii) are not equivalent and in fact are off by 1. We also know (i) is equivalent to our asked expression, allowing us to eliminate options (B) and (D).

OpenStudy (anonymous):

why isn't (i) right?

OpenStudy (anonymous):

It is a prereq class. I have to take 2 of these before I can start any college level math courses.

OpenStudy (anonymous):

jesus they must hate you

OpenStudy (anonymous):

oh i see @oldrin.bataku you did not say it was not right, sorry

OpenStudy (anonymous):

since \(3333+3333=6666\) base 7, it cannot also be \(1000\) base seven as well that is for sure, so scratch (iii)

OpenStudy (anonymous):

Lol, I think they do. Oh, the best part is I still pay for them but it doesn't count as any credits.

OpenStudy (anonymous):

i think both (i) and (ii) are right

OpenStudy (anonymous):

they are both 2400 base ten

OpenStudy (anonymous):

as a matter of fact i am sure of it

OpenStudy (anonymous):

oops missed the fact there's still a \(7^0\) there!

OpenStudy (anonymous):

go with C (i) and (ii) are true

OpenStudy (anonymous):

Thanks everyone

OpenStudy (anonymous):

yw

jimthompson5910 (jim_thompson5910):

6666 base 7 = 6*7^3 + 6*7^2 + 6*7^1 + 6*7^0 base 10 6666 base 7 = 2400 base 10 -------------------- now convert to base 12 2400/12 = 200 remainder 0 200/12 = 16 remainder 8 16/12 = 1 remainder 4 1/12 = 0 remainder 1 read the remainders from bottom to top to get 1480 ------------------- so this means 6666 base 7 = 2400 base 10 = 1480 base 12